OpenStudy (anonymous):

I NEED HELP...NOW I CANNOT GET THIS! A cliff diver dives off a cliff 40 feet above water. Write an equation giving the diver's height H (in feet) above the water after t seconds. How long is the diver in the air?

OpenStudy (anonymous):

\[ H(t) = -16t^2 + V_0 t + H_0\] In your situation, \[V_0 = 0\text{, and }H_0 = 40\] You can use the quadratic equation to solve for t. If t < 0 it doesn't make any sense in this situation so just throw it out and stick with the positive value(s).

OpenStudy (anonymous):

1. h = 40 - 16t^2 0 = 40 - 16t^2 16t^2 = 40 t^2 = 5/2 t = √(5/2) ≈ 1.58 sec

OpenStudy (anonymous):

where did you get -16 from?

OpenStudy (anonymous):

that's the acceleration due to the earth's gravity (in ft/(sec^2))

OpenStudy (anonymous):

THANKS!

OpenStudy (anonymous):

you're welcome!